# The Open Journal of Astrophysics

### v1

@@ 3.5 Relating to established boundary definitions
Figure 8: Dependence of the transition region on host mass using 1/1 labels based on three established cluster boundaries ($R_{\mathrm{sp}}$, $R_{200,\mathrm{mean}}$, and $R_{200,\mathrm{crit}}$) with six primary properties. This is overlaid with the same reference curves from Diemer (2020) and More, Diemer & Kravtsov (2015). This indicates that these boundary definitions do not distinctly separate cluster and field galaxies based on their intrinsic properties.
@@ 3.6 Property‑based transition boundaries
Figure 9: Median zero-crossing radius ratios between different physical galaxy property categories. The top row uses the fiducial distance-based $0.5/5.0$ label. The bottom row employs the SubLink label. Each panel shows the ratio of transition radii measured from one property category to another, plotted as a function of cluster mass $M_{200,\mathrm{mean}}$. Line legends on the panels denote whether sSFR is counted a stellar or gas property. For each mass bin, the median ratio among these models is computed. This shows that there exist offsets between where different property categories experience transition, indicating that the underlying physical processes can be inferred spatially.
@@ 4 Discussion of limitations and future work
The central question addressed here is: Can galaxy membership probabilistically inform the spatial extent of a cluster’s environmental influence? Future work could be proposed to resolve the reciprocal problem: Can a boundary inferred purely from galaxy properties be used to determine cluster membership? One possible direction is to employ unsupervised clustering techniques within the high‐dimensional galaxy property space to identify spatial separations that may correspond to physical boundaries. Such methods could yield a more objective, data‐driven spatial demarcation between cluster and field populations, enabling a direct comparison between spatially and property‐based definitions of cluster extent.
Beyond the mass-based cluster environment selection criteria done in this work, we can also choose cluster environments based on whether they host AGNs. This allows us to further probe black hole related quenching mechanisms and investigate the spatial effect of AGNs on galaxy intrinsic properties.
This framework can also be applied to observations. Finding a transition boundary in cluster environments solely based on observables can illuminate observational work. Redshift evolution is also worth investigating since cluster environments evolve significantly with time (O’Neil et al., 2024). Comparisons with other simulations can be conducted to reveal more information.
@@ 5 Conclusions
  • In Figure 3, we provide the first quantitative, probabilistic map of the transition region, which directly links galaxy–cluster distance to membership probability. From this result, we further extract the cluster-to-cluster zero point median and 16th-84th percentile to further highlight the extent of the cluster environment’s influence as shown in Figure 5. We find that the transition region determined probabilistically using our algorithm follows an increasing trend with the cluster environment mass. This region is intrinsically scattered, highlighting the dynamical trajectory of galaxies entering cluster environments. The mass dependence of the transition region is consistent with the place ram-pressure stripping occurs for high-mass clusters compare to low-mass ones. Figure 7 shows that the mass dependence follows the power law $x^{0.1}$.
@@
The probabilistic nature of the transition boundary in cluster outskirts is a signature of mixed populations. A possible future test would be to utilise kinematic data from upcoming surveys like DESI or Roman. By first separating galaxies into infalling and orbiting samples, one could apply our classifier to each population independently. This would directly test whether the intrinsic scattering is resolved and reveal if each population has its own, sharper transition profile.
@@ 6 Acknowledgements
We thank Daisuke Nagai, Bhuvnesh Jain, Tri Nguyen, Frank van den Bosch, and Jeffrey Kenney for helpful discussions. C.H. was supported by the Pamela Daniels’59 Fellowship at Wellesley College, the Wellesley College Career Education Grant, and the MIT Undergraduate Research Opportunities Program (UROP). S.O. is supported by the National Science Foundation under Grant No. AST-2307787. Some of the computations were performed on the Engaging cluster supported by the Massachusetts Institute of Technology.

### v2

@@ 3.5 Relating to established boundary definitions
Figure 8: Dependence of the transition region on host mass using 1/1 labels based on three established cluster boundaries ($R_{\mathrm{sp}}$, $R_{200,\mathrm{mean}}$, and $R_{200,\mathrm{crit}}$) with six primary properties. This is overlaid with the same reference curves from Diemer (2020) and More, Diemer & Kravtsov (2015). If these boundaries are good demarcations to determine cluster and field galaxies, then a 1-to-1 label scheme should result in a clean separation on this end, where the transition point for each mass bin equals to the value of the particular boundary of clusters with the same mass. There also should be no error bars. This indicates that these boundary definitions do not distinctly separate cluster and field galaxies based on their intrinsic properties.
@@ 3.6 Property‑based transition boundaries
Figure 9: Median ratios of zero-crossing (transition) radii measured from different property categories. The top row uses the fiducial distance-based $0.5/5.0$ label, and the bottom row uses the SubLink label. Within each row, the columns correspond to different radius ratios, where the leftmost column is $R_{\star}/R_{\rm gas}$, middle column is $R_{\star}/R_{\rm dyn}$, and the rightmost column is $R_{\rm gas}/R_{\rm dyn}$. The variables $R_{\star}$, $R_{\rm gas}$, and $R_{\rm dyn}$ denote the transition radius inferred using the stellar-, gas-, and dynamical-intrinsic galaxy-property feature sets, respectively. The ratios are plotted as a function of cluster mass $M_{200,\mathrm{mean}}$. For each mass bin we take the median ratio across clusters (connected by lines). The horizontal dashed line marks when the the two property categories have identical transition radii. Because sSFR can be treated as either a stellar- or gas-related feature, we show several model variants as denoted by the legend labels. In the stellar/gas ratio panels, four curves indicate whether sSFR is excluded from both sets, included in both sets, or included in only one set. In the stellar/dynamical and gas/dynamical panels, two curves indicate whether sSFR is included or excluded from the relevant baryonic property set. Overall, deviations from unity indicate offsets between the radii at which different property categories transition, consistent with different physical processes acting over different spatial and temporal scales.
@@ 4 Discussion of limitations and future work
The central question addressed here is: Can galaxy membership probabilistically inform the spatial extent of a cluster’s environmental influence? Future work could be proposed to resolve the reciprocal problem: Can a boundary inferred purely from galaxy properties be used to determine cluster membership? One possible direction is to employ unsupervised clustering techniques within the high‐dimensional galaxy property space to identify spatial separations that may correspond to physical boundaries. Such methods could yield a more objective, data‐driven spatial demarcation between cluster and field populations, enabling a direct comparison between spatially and property‐based definitions of cluster extent. It is important to note that the properties incorporated in this study are difficult to disentangle.
Beyond the mass-based cluster environment selection criteria done in this work, we can also choose cluster environments based on whether they host AGNs (Wing & Blanton, 2011; Galametz et al., 2012; Wylezalek et al., 2013; Croston et al., 2019; Golden-Marx et al., 2019). This allows us to further probe black hole related quenching mechanisms and investigate the spatial effect of AGNs on galaxy intrinsic properties.
This framework can also be applied to observations. During its observational adaptation, the training and validation sets of the DNN can be created with the help of cluster finders like the the redMaPPer (Rykoff et al., 2014, 2016). The resulting probabilistic transition region can also be statistically weighted with the redMaPPer’s outcome. Several observable properties can be chosen in this direction including multi-band colours, magnitudes, surface brightness, and spectral lines with photometric and redshift information. Projection effects (line-of-sight superpositions and photometric-redshift uncertainties) can lead to interloper contamination in galaxy membership assignments, which boosts the inferred richness and increases the scatter in any inferred transition radius. Therefore, the probabilistic transition region will be broadened due to potential biases compared to the intrinsic 3D case (Rozo et al., 2015; Sunayama et al., 2020; Costanzi et al., 2018; Myles et al., 2025).
Redshift evolution is also worth investigating since cluster environments evolve significantly with time (O’Neil et al., 2024). Comparisons with other simulations can be conducted to reveal more information, especially by adding dwarf galaxies to illuminate stronger implications of the cluster environment’s impact on low-mass systems.
@@ 5 Conclusions
  • In Figure 3, we provide the first quantitative, probabilistic map of the transition region, which directly links galaxy–cluster distance to membership probability. In Figure 5, we further extract the cluster-to-cluster zero point median and with its the 16th-84th percentile errors to highlight the transition region marked by the cluster environment’s influence on galaxies. It follows an increasing trend with the cluster environment mass, which is consistent with the location of ram pressure stripping for clusters of different masses. Figure 7 shows that the mass dependence follows the power law $x^{0.1}$. It is also intrinsically scattered, which means that the process of a galaxy’s properties changing when entering the cluster environment is gradual.
@@
The probabilistic nature of the transition boundary in cluster outskirts is a signature of mixed populations. A possible future test would be to utilise kinematic data from ongoings surveys such as DESI (DESI Collaboration et al., 2022, 2024a, 2024b, 2025) or upcoming surveys such as Roman (Spergel et al., 2015). By first separating galaxies into infalling and orbiting samples, one could apply our classifier to each population independently. This would directly test whether the intrinsic scattering is resolved and reveal if each population has its own, sharper transition profile.
@@ 6 Acknowledgements
We thank the Editor and the Referee for their careful reading and constructive comments. We thank Daisuke Nagai, Bhuvnesh Jain, Tri Nguyen, Lina Necib, Frank van den Bosch, and Jeffrey Kenney for helpful discussions. C.H. was supported by the Pamela Daniels’59 Fellowship at Wellesley College, the Wellesley College Career Education Grant, and the MIT Undergraduate Research Opportunities Program (UROP). S.O. is supported by the National Science Foundation under Grant No. AST-2307787. Some of the computations were performed on the Engaging cluster supported by the Massachusetts Institute of Technology.